# Spectral Super-Resolution Neural Operator with Atmospheric Radiative Transfer Prior

### v1

@@ Abstract
Spectral super-resolution (SSR) aims to reconstruct hyperspectral images (HSIs) from multispectral observations, with broad applications in remote sensing. Data-driven methods are widely used, but they often overlook physical principles, leading to unrealistic spectra, particularly in atmosphere-affected bands. To address this challenge, we propose the Spectral Super-Resolution Neural Operator (SSRNO), which incorporates atmospheric radiative transfer (ART) prior into the data-driven procedure, yielding more physically consistent predictions. The proposed SSRNO framework consists of three stages: upsampling, reconstruction, and refinement. In the upsampling stage, we leverage prior information to expand the input multispectral image, producing a physically plausible hyperspectral estimate. Subsequently, we utilize a neural operator in the reconstruction stage to learn a continuous mapping across the spectral domain. Finally, the refinement stage imposes a hard constraint on the output HSI to eliminate color distortion. The upsampling and refinement stages are implemented via the proposed guidance matrix projection (GMP) method, and the reconstruction neural operator adopts U-shaped spectral-aware convolution (SAC) layers to capture multi-scale features. Moreover, we theoretically demonstrate the optimality of the GMP method. With the neural operator and ART priors, SSRNO also achieves continuous spectral reconstruction and zero-shot extrapolation. Various experiments validate the effectiveness and generalization ability of the proposed approach.
@@ I Introduction
Traditional SSR approaches mainly include early model-based methods [29], regression-based methods [31, 41], and sparse representation methods [1, 49]. In remote sensing, they have been applied in conjunction with domain-specific priors, such as spectral unmixing features [47]. Traditional methods primarily rely on prior knowledge and physical constraints, which contribute to robustness and interpretability. However, they often fail to fully leverage large-scale data, which may lead to underfitting.
@@
Advances in deep learning have led to a range of approaches based on deep neural networks. Numerous methods employing convolutional neural networks (CNNs) [37, 30] and attention mechanisms [3, 42, 20] lead to notable improvements and outperform traditional approaches. For remote sensing applications, deep learning-based methods are also widely adopted. For example, fusion-based methods [32, 9] use paired low-resolution HSIs and high-resolution MSIs to extract and fuse spectral-spatial information, producing high-resolution hyperspectral images. However, the practical application of fusion-based methods is hindered by the difficulty of obtaining registered HSI-MSI pairs. Thus, single-image spectral super-resolution methods [8, 26, 20, 48, 51] have gained attention, as they require only the input MSI or RGB image at the inference stage. For example, Zhou et al. [51] proposed a CNN-Transformer hybrid network for directly reconstructing hyperspectral images from RGB inputs, thereby enhancing spectral super-resolution. Unsupervised spectral super-resolution has also been extensively studied to address the challenge of acquiring paired MSI-HSI data [34, 7, 21]. Additionally, several studies aim to recover hyperspectral spectra at arbitrary scale [42, 19, 43], drawing inspiration from the success of implicit neural representations [5], which have been widely used in spatial super-resolution.
Data-driven methods achieve strong performance, but they often overlook underlying physical principles, resulting in unrealistic predictions. To address this, some works have sought to embed physics or priors into the data-driven framework to achieve both physical interpretability and high performance [24, 13, 33, 15, 16, 40, 21]. For example, Lin et al. [24] incorporated spectral response functions into CNN-based models to achieve physically plausible spectral recovery. Wu et al. [40] embedded comprehensive priors to regularize and optimize the solution space of SSR. Huo et al. [16] integrated the physical relationship between spectral reflectance and corresponding RGB images via a meta-auxiliary learning strategy.
While physics-informed methods improve the plausibility of spectral reconstruction, many neglect key remote sensing factors such as solar irradiance and atmospheric transmittance, resulting in discrepancies with actual image formation. In addition, many methods are limited to a fixed number of spectral bands, restricting their applicability for arbitrary or high-resolution spectral outputs.
@@
To address these challenges, we propose the Spectral Super-Resolution Neural Operator (SSRNO), which embeds extraterrestrial irradiance and direct beam radiation predicted by atmospheric radiative transfer (ART) codes into a data-driven framework. Inspired by the two-stage paradigm [37], the proposed framework comprises two standard stages, upsampling and reconstruction, and an additional refinement stage, as illustrated in Figure 2. In the upsampling stage, the spectral response function (SRF) and ART priors are leveraged to generate an image with higher spectral resolution. With the upsampled HSI, we employ a data-driven neural operator to reconstruct a finer HSI estimate. Since the neural operator is resolution-invariant and coordinate-based [23, 18], we can compute the estimated HSI at arbitrary spectral scales. Finally, we refine the HSI by projecting it onto the affine space defined by the SRF, thereby ensuring consistency with the original MSI after degradation. The upsampling and refinement stages employ our proposed guidance matrix projection (GMP), a novel methodology derived from an affine space projection problem. It not only leverages physical priors to upsample the MSI but also enforces hard constraints to guarantee reconstruction consistency. Additionally, we employ a reconstruction neural operator with spectral-aware convolution (SAC) layers in a U-shaped structure to capture multi-scale spatial and spectral features. With the ART prior and the neural operator, SSRNO achieves superior continuous and zero-shot reconstruction capabilities.
@@
  • We propose a physics-informed approach for spectral super-resolution, which integrates ART priors into the data-driven framework to produce physically plausible predictions;
  • We derive a novel and versatile method from an affine projection formulation, enabling both upsampling and constraint incorporation, with theoretical guarantees of optimality;
  • We apply the neural operator architecture to spectral reconstruction, providing physically grounded interpretability and supporting continuous spectral reconstruction.
II Preliminaries
In this section, we introduce the preliminary works upon which our approach is built. We first introduce the atmospheric radiative transfer model in Section II-A, which forms the foundation of the physical prior used in our method. We then present the operator learning framework in Section II-B, which motivates the design of our network architecture.
II-A Atmospheric Radiative Transfer Model
Figure 1: Comparison to the extraterrestrial irradiance $E_{on\lambda}$ and the direct beam radiation $E_{bn\lambda}$, where $E_{on\lambda}$ is based on the Air Mass 1.5 spectra and $E_{bn\lambda}$ is based on the 1976 U.S. Standard Atmosphere.
Solar radiation and atmospheric transmission modeling are crucial components of many scientific disciplines. Over the past several decades, atmospheric radiative transfer models have undergone significant evolution to meet the demands of applications ranging from climate modeling to agriculture and remote sensing. Generally, these models aim to maintain highly accurate irradiance predictions under specific circumstances, such as vertical atmospheric profiles, seasonal variations, and geographical latitude. Examples of such models include MODerate resolution atmospheric TRANsmission (MODTRAN) [2] and Santa Barbara Disort Atmospheric Radiative Transfer (SBDART) [36]. Moreover, the Simple Model of the Atmospheric Radiative Transfer of Sunshine (SMARTS) [11] is developed to overcome the limitations of its predecessor by offering significantly improved spectral resolution, expanded atmospheric constituent modeling, and bridging the gap between high-accuracy scientific models and computational efficiency needed for solar engineering [12]. Estimating a wide spectral range from 280 nm to 4000 nm, the SMARTS includes parameterizations of gas absorption, aerosol scattering, and spectral albedo, and has been instrumental in generating standard solar spectra for photovoltaic (PV) system testing. In this article, the ART prior is from the SMARTS.
In SMARTS, the beam irradiance received at ground level by a surface oriented perpendicular to the sun rays at wavelength $\lambda$ is defined as
$E_{bn\lambda}=E_{on\lambda}T_{R\lambda}T_{o\lambda}T_{n\lambda}T_{g\lambda}T_{w\lambda}T_{a\lambda},$ (1)
where $E_{bn\lambda}$ is the direct beam radiation, $E_{on\lambda}$ is the extraterrestrial irradiance, and the transmittance factors $T_{*\lambda}$ represent various atmospheric effects as summarized in Table I. The transmittance factors, as detailed in [11], can be computed based on the given geophysical conditions. As shown in Figure 1, the direct beam radiation exhibits several distinct absorption bands compared to the reference extraterrestrial irradiance, primarily in the range of 750-2000 nm. These features are mainly attributed to strong absorption by water vapor (H2O), carbon dioxide (CO2), and oxygen (O2). Besides, the spectral irradiance below 400 nm is also significantly absorbed by ozone (O3). In the spectral reconstruction of remote sensing imagery, the atmosphere has a substantial impact on specific wavelength bands, making it a physical factor that cannot be overlooked.
TABLE I: Atmospheric Transmittance Factors of SMARTS
Factors Extinction process
$T_{R\lambda}$ Rayleigh scattering
$T_{o\lambda}$ Ozone ($\text{O}_{3}$) absorption
$T_{n\lambda}$ Nitrogen dioxide ($\text{NO}_{2}$) absorption
$T_{g\lambda}$ Uniformly mixed gas (mainly $\text{O}_{2}$, $\text{CO}_{2}$) absorption
$T_{w\lambda}$ Water vapor ($\text{H}_{2}\text{O}$) absorption
$T_{a\lambda}$ Aerosol extinction
II-B Operator Learning
Figure 2: Overall framework of the proposed method. The input MSI undergoes three main stages to output the final HSI $\hat{Y}$: (1) Upsampling, where we interpolate and extrapolate the input MSI to the upsampled HSI, (2) Reconstruction, where we employ the neural operator to transform the upsampled HSI to a finer HSI estimate, and (3) Refinement, where we impose a hard constraint on the final result to enforce zero reconstruction error and eliminate color distortion.
@@ II-B Operator Learning
Operator learning is an emerging framework in machine learning that extends the learning task from finite-dimensional vector mappings to mappings between infinite-dimensional function spaces. Lu et al. [27] first established the deep operator network architecture, which is capable of learning continuous mappings and supporting queries at arbitrary scales.
@@
Recently, a new model termed the neural operator has been proposed, which involves using neural networks to approximate mesh-free, infinite-dimensional, and nonlinear operators [18, 23, 35]. Let $a:D\to\mathbb{R}^{d_{a}}$ denote the input function defined on a bounded domain $D\subset\mathbb{R}^{d}$, $u:D^{\prime}\to\mathbb{R}^{d_{u}}$ denote the output function defined on $D^{\prime}\subset\mathbb{R}^{d^{\prime}}$. Neural operators $\mathcal{G}_{\theta}$ learn the non-linear mapping $\mathcal{G}^{{\dagger}}$ between two infinite-dimensional spaces by using a finite collection of observations of input-output pairs from this mapping, which is,
@@ III Methodology
Figure 3: Overview of the upsampling, reconstruction, and refinement stages. The upsampling and refinement stages both employ the GMP method, but with different objectives. In the upsampling stage, GMP incorporates prior knowledge to produce a physically plausible HSI estimate, whereas in the refinement stage, it projects the HSI onto the affine space defined by (4). The reconstruction stage employs a neural operator with a U-shaped structure and spectral-aware convolution layers to learn the mapping from upsampled HSIs to target HSIs across continuous spectral dimensions.
@@
In this section, we present the SSRNO methodology. We begin by formulating the spectral reconstruction problem and outlining the motivation and overall framework of our approach in Section III-A. We then introduce two key components of the framework: the proposed guidance matrix projection method and the U-shaped reconstruction neural operator with spectral-aware kernel integrals. Section III-B discusses the GMP method and its theoretical optimality, while Section III-C provides a detailed description of the neural operator architecture.
III-A Motivation and Overall Framework
@@
Let $Y\in\mathbb{R}^{C\times N}$ denote the HSI, $X\in\mathbb{R}^{M\times N}$ the paired RGB or multispectral image, and $S\in\mathbb{R}^{M\times C}$ the SRF, where $N=HW$ is the number of pixels, and $C$ and $M$ are the number of channels with $C>M$. Generally, SRFs contain enough information for each spectral band, so we assume that all SRF matrices are of full row rank. The relation between $X$ and $Y$ can be mathematically formulated in a linear matrix equation, that is,
$SY=X.$ (4)
@@
Thus, the spectral reconstruction problem can be formally formulated as solving the underdetermined linear matrix equation (4), where, given an RGB or multispectral image $X$ and its corresponding spectral response function $S$, the objective is to recover a plausible solution $\hat{Y}$ to (4).
@@
To solve this ill-posed problem, a widely used approach is the two-stage reconstruction paradigm, where the input is first upsampled to the target resolution or modality space, followed by a data-driven reconstruction module that restores missing details and corrects estimation errors [37]. This paradigm can be formulated as
$\hat{Y}=R(U(X))+U(X),$ (5)
@@
where $U(\cdot)$ denotes the upsampling operation that projects the input $X$ into a higher-dimensional target space, and $R(\cdot)$ is a learnable data-driven reconstruction module, such as CNN-based neural networks. A residual connection is employed to facilitate learning and enhance reconstruction quality.
@@
However, this paradigm faces inherent limitations when applied to spectral reconstruction. For the upsampler $U(\cdot)$, commonly used techniques such as bilinear interpolation and transposed convolution are less effective in spectral super-resolution, as the upsampling ratio is substantially higher than in typical spatial super-resolution tasks. Additionally, this paradigm is entirely data-driven, lacking integration of physical processes such as atmospheric transmission and sensor response that are essential for spectral reconstruction. Besides, most data-driven reconstruction modules rely on deep neural networks that learn mappings between finite-dimensional spaces. However, such designs neglect the continuity of spectral signatures, thereby limiting both performance and applicability.
@@
Hence, inspired by (5) and prior studies on multistage frameworks [22], we adopt a similar yet more effective three-stage design, as illustrated in Figure 2. Given the MSI $X$, wavelength coordinates $\boldsymbol{w}$, and the SRF $S_{\boldsymbol{w}}$ discretized at $\boldsymbol{w}$, we sequentially apply the following three stages to obtain the final HSI estimate.
@@
    Upsampling: In the first upsampling stage, we apply the proposed GMP method to obtain the initial upsampled HSI, which can be formulated as
    $\bar{Y}=U(X|Z_{\boldsymbol{w}},S_{\boldsymbol{w}})=P_{\text{GMP}}(Z_{\boldsymbol{w}},S_{\boldsymbol{w}},X),$ (6)
    where $U(\cdot|Z_{\boldsymbol{w}},S_{\boldsymbol{w}})$ is the conditional upsampler guided by $Z_{\boldsymbol{w}}$ and $S_{\boldsymbol{w}}$, $P_{\text{GMP}}(\cdot,\cdot,\cdot)$ is our proposed guidance matrix projector with given pixel-wise prior $Z_{\boldsymbol{w}}\in\mathbb{R}^{C\times N}$ as the guidance matrix, SRF $S_{\boldsymbol{w}}\in\mathbb{R}^{M\times C}$ and MSI $X\in\mathbb{R}^{M\times N}$, which are all discretized at $\boldsymbol{w}$.
@@
    $\tilde{Y}=R_{no}(\bar{Y},\boldsymbol{w})+\bar{Y},$ (7)
    where $\tilde{Y}$ is the reconstructed HSI, $R_{no}(\cdot,\cdot)$ is the neural operator that takes the functional value and the coordinate as inputs, and a residual connection is employed similar to Equation 5.
@@
    Refinement: In the final refinement stage, we project our final output to the solution space of (4), which imposes a hard constraint on our result. This stage can be formulated as
    $\hat{Y}=R_{f}(\tilde{Y}|S_{\boldsymbol{w}},X)=P_{\text{GMP}}(\tilde{Y},S_{\boldsymbol{w}},X),$ (8)
    where $\hat{Y}$ is the output HSI, $R_{f}(\cdot|S_{\boldsymbol{w}},X)$ is the refinement operator. Although the $R_{f}(\cdot|S_{\boldsymbol{w}},X)$ employs the same GMP procedure as the upsampler in the first stage, its objective is fundamentally different. The upsampling stage aims to produce a physically plausible upsampled HSI. In contrast, the refinement stage projects our final output to the solution space of (4), which imposes a hard constraint on our result.
Our framework successfully addresses the limitations of the traditional two-stage paradigm. First, the GMP method embeds physical priors into the data-driven framework through pixel-wise guidance. It also refines the output by solving a constrained optimization problem, thereby enforcing the hard constraint defined in (4). Moreover, previous studies have shown that neural operators can accommodate inputs defined on meshes with different discretizations [23, 18], so we employ a neural operator to address the inherent limitation of conventional neural networks that can only model discrete spectral vectors.
Within the proposed framework, the guidance matrix projection method and the reconstruction neural operator serve as two core components, which are detailed in Section III-B and Section III-C, respectively.
III-B Guidance Matrix Projection
In this section, we propose a novel method termed Guidance Matrix Projection, based on the formulation in Section III-A. We seek a physically plausible HSI estimate that remains close to the guidance matrix $Z$ while satisfying the constraint in (4). To achieve this, we can consider the following affine space projection problem:
$\min_{Y}$ $\quad\text{SAM}(Y,Z)$ (9)
s.t. $\quad SY=X,$

### v2

# Radiative-Structured Neural Operator for Continuous and Extrapolative Spectral Super-Resolution

@@ Abstract
Spectral super-resolution (SSR) aims to reconstruct hyperspectral images (HSIs) from multispectral observations, with broad applications in computer vision and remote sensing. Deep learning-based methods have been widely used, but they often treat spectra as discrete vectors learned from data, rather than continuous curves constrained by physics principles, leading to unrealistic predictions and limited applicability. To address this challenge, we propose the Radiative-Structured Neural Operator (RSNO), which learns a continuous mapping for spectral super-resolution while enforcing physical consistency under the radiative prior. The proposed RSNO consists of three stages: upsampling, reconstruction, and refinement. In the upsampling stage, we leverage prior information to expand the input multispectral image, producing a physically plausible hyperspectral estimate. Subsequently, we adopt a neural operator backbone in the reconstruction stage to learn a continuous mapping across the spectral domain. Finally, the refinement stage imposes a hard constraint on the output HSI to eliminate color distortion. The upsampling and refinement stages are implemented via the proposed angular-consistent projection (ACP), which is derived from a non-convex optimization problem. Moreover, we theoretically demonstrated the optimality of ACP by null-space decomposition. Various experiments validate the effectiveness of the proposed approach across conventional spectral super-resolution, continuous spectral reconstruction, and infrared extrapolation.
@@ I Introduction
Traditional SSR approaches mainly include early model-based methods [29], regression-based methods [33, 44], and sparse representation methods [1, 49]. These methods primarily rely on prior knowledge and physical constraints, which contribute to robustness and interpretability. However, they often fail to fully leverage large-scale data, which may lead to underfitting.
@@
Advances in deep learning have led to a range of approaches based on deep neural networks. Numerous methods employing convolutional neural networks (CNNs) [38, 30] and attention mechanisms [3, 45, 20] lead to notable improvements and outperform traditional approaches. Deep learning-based methods can also be categorized according to their problem settings. For example, fusion-based methods [34, 10] use paired low-resolution HSIs and high-resolution MSIs to extract and fuse spectral-spatial information, producing high-resolution hyperspectral images. However, the practical application of fusion-based methods is hindered by the difficulty of obtaining registered HSI-MSI pairs. Thus, single-image spectral super-resolution methods [41, 26, 20, 48, 51] have gained attention, as they require only the input MSI or RGB image at the inference stage. For example, Zhou et al. [51] proposed a CNN-Transformer hybrid network for directly reconstructing hyperspectral images from RGB inputs, thereby enhancing spectral super-resolution. Unsupervised spectral super-resolution has also been extensively studied to address the challenge of acquiring paired MSI-HSI data [35, 8].
Deep learning-based approaches have achieved great success, but they typically formulate SSR as a finite-dimensional regression task learned purely from data. In reality, spectra are continuous functions shaped by inherent physical priors, rather than discrete vectors produced by data-driven models. Ignoring this fundamental property may result in physically inconsistent predictions, degraded generalization ability, and limited practical applicability.
@@
To address this challenge, we propose the Radiative-Structured Neural Operator (RSNO), which models spectra in infinite-dimensional space and incorporates physically grounded spectral priors derived from atmospheric radiative transfer models. Inspired by the two-stage paradigm [38, 30, 21], the proposed framework comprises two standard stages, upsampling and reconstruction, and an additional refinement stage, as illustrated in Figure 1. In the upsampling stage, the spectral response function (SRF) and radiative transfer priors are leveraged to generate an image with higher spectral resolution. With the upsampled HSI, we employ a data-driven neural operator backbone to reconstruct a finer HSI estimate. Since the neural operator is resolution-invariant and coordinate-based [22, 18], we can compute the estimated HSI at arbitrary spectral scales. Finally, we refine the HSI by projecting it onto the affine space defined by the SRF, thereby ensuring consistency with the original MSI after degradation. The upsampling and refinement stages employ our proposed angular-consistent projection (ACP), a novel methodology derived from a non-convex affine space projection problem based on cosine similarity. It not only leverages physical priors to upsample the MSI but also enforces hard constraints to guarantee reconstruction consistency. Additionally, the neural operator backbone adopts a U-shaped architecture equipped with spectral-aware convolution (SAC) layers to capture multi-scale spatial–spectral features. Experiments on real-world datasets show that RSNO not only performs competitively in conventional finite-dimensional SSR, but also enables continuous spectral reconstruction and infrared extrapolation, highlighting its effectiveness and flexibility in practical applications.
@@
  • We incorporate atmospheric radiative transfer priors into the data-driven framework, enabling radiative-structured and physically coherent spectral reconstruction;
  • We formulate spectral super-resolution as a continuous regression problem in infinite-dimensional space, supporting both continuous interpolation and infrared extrapolation;
  • We derive the ACP method from a non-convex affine space projection problem based on cosine similarity, and theoretically demonstrate its optimality by null-space decomposition.
II Related work
II-A Continuous Spectral Super-Resolution
Continuous spectral super-resolution aims at recovering spectral curves at arbitrary resolutions. Xu et al. [45] first introduced a framework that reconstructs spectral images with an arbitrary number of bands and demonstrated superior performance over discrete methods in fixed-band reconstruction tasks. Beyond this, they further upgraded the neural attention mapping module in their subsequent work [46], thereby enabling continuous reconstruction in both the spatial and spectral domains with state-of-the-art performance among discrete baselines. For HSI reconstruction of the Coded Aperture Snapshot Spectral Imaging (CASSI) system, Chen et al. [5] proposed a spectral-wise attention mechanism with structural designs to achieve continuous HSI recovery. Similarly, Li et al. [19] proposed a lightweight architecture that supports continuous reconstruction across both spatial and spectral dimensions within the CASSI framework.
These approaches are all built upon implicit neural representations (INRs) [6]. Originating from the neural radiance field (NeRF) [31], INR implicitly represents natural and complex images continuously and has inspired a wide range of arbitrary-scale super-resolution methods in computer vision [32].
II-B Physics-Informed Spectral Super-Resolution
While data-driven methods have achieved strong performance in spectral super-resolution, their lack of explicit physical constraints has motivated the development of physics-informed approaches [23, 13, 25, 16, 43, 7]. These methods seek to integrate physical models or domain-specific priors into learning-based frameworks to enhance reconstruction fidelity and interpretability.
Existing physics-informed SSR methods can be broadly categorized into regularization-based and model-embedded approaches, depending on how physical priors are incorporated. Regularization-based methods typically incorporate physical knowledge at the objective level in the form of auxiliary losses or regularization terms. For example, Wu et al. [43] embedded comprehensive semantic priors to regularize and optimize the solution space of SSR. Model-embedded methods integrate physical knowledge directly into the model design, where physical structures are reflected in the network architecture, parameterization, or computational modules. For example, Liu et al. [25] embedded a spectral unmixing model into the reconstruction framework and explicitly accounted for solar radiation and atmospheric absorption effects. In addition, several works [23, 16, 8] approached SSR from the perspective of a linear inverse problem and derived null-space priors to constrain the solution space, realizing high-fidelity reconstruction.
II-C Operator Learning
@@
Operator learning is an emerging framework in machine learning that extends the learning task from finite-dimensional vector mappings to mappings between infinite-dimensional function spaces. Representative architectures include DeepONet [27] and neural operator [22, 18, 36].
@@
The neural operator is formulated as follows. Let $a:D\to\mathbb{R}^{d_{a}}$ denote the input function defined on a bounded domain $D\subset\mathbb{R}^{d}$, $u:D^{\prime}\to\mathbb{R}^{d_{u}}$ denote the output function defined on $D^{\prime}\subset\mathbb{R}^{d^{\prime}}$. Neural operators $\mathcal{G}_{\theta}$ learn the non-linear mapping $\mathcal{G}^{{\dagger}}$ between two infinite-dimensional spaces by using a finite collection of observations of input-output pairs from this mapping, which is,
@@
II-D Atmospheric Radiative Transfer Model
Atmospheric radiative transfer models are crucial components of many scientific disciplines. Generally, these models aim to estimate atmospheric effects on the solar radiation under specific circumstances, such as vertical profiles, seasonal variations, and geographical latitude. Examples of atmospheric radiative transfer models mainly include MODerate resolution atmospheric TRANsmission (MODTRAN) [2], Santa Barbara Disort Atmospheric Radiative Transfer (SBDART) [37], and the Simple Model of the Atmospheric Radiative Transfer of Sunshine (SMARTS) [11, 12]. In this article, the radiative spectral prior is predicted by the SMARTS.
@@ III Methodology
Figure 1: Overall framework of the proposed method. The input MSI undergoes three main stages to output the final HSI $\hat{X}$: (1) Upsampling, where we expand the input MSI to the upsampled HSI, (2) Reconstruction, where we employ the neural operator to transform the upsampled HSI to a finer HSI estimate, and (3) Refinement, where we impose a hard constraint on the final result to enforce zero reconstruction error and eliminate color distortion. The upsampling and refinement stages both employ the ACP method, but with different objectives. In the upsampling stage, ACP incorporates radiative spectral prior to produce a radiative-structured and physically plausible HSI estimate, whereas in the refinement stage, it projects the HSI onto the affine space defined by (4).
@@
In this section, we present the RSNO methodology. We begin by formulating the spectral reconstruction problem and outlining the motivation and overall framework of our approach in Section III-A. We then introduce two key components of the framework: the proposed angular-consistent projection method and the U-shaped reconstruction neural operator with spectral-aware kernel integrals. Section III-B discusses the ACP method and its theoretical optimality, while Section III-C provides a detailed description of the neural operator backbone.
III-A Overall Framework
@@
Let $X\in\mathbb{R}^{C\times N}$ denote the HSI, $Y\in\mathbb{R}^{M\times N}$ the paired RGB or multispectral image, and $S\in\mathbb{R}^{M\times C}$ the SRF, where $N=HW$ is the number of pixels, and $C$ and $M$ are the number of channels with $C>M$. Generally, SRFs contain enough information for each spectral band, so we assume that all SRF matrices are of full row rank. The forward degradation process from the HSI $X$ to the MSI $Y$ can be mathematically formulated in a linear inverse problem,
$Y=SX+\mathcal{E},$ (3)
where $\mathcal{E}\in\mathbb{R}^{M\times N}$ is the pixel-wise noise. To statistically fit the model, we omit the noise and simplify it to the following underdetermined linear matrix equation:
$Y=SX.$ (4)
@@
Thus, the spectral reconstruction problem can be formally formulated as solving (4), where, given an RGB or multispectral image $Y$ and its corresponding spectral response function $S$, the objective is to recover a plausible solution $\hat{X}$ to (4).
@@
To solve this ill-posed problem, a widely used deep learning-based approach is the two-stage reconstruction paradigm, where the input is first upsampled to the target resolution, followed by a reconstruction neural network that restores missing details and corrects estimation errors [38, 30, 21]. This paradigm can be formulated as
$\hat{X}=R(U(Y))+U(Y),$ (5)
@@
where $U(\cdot)$ denotes the upsampling operation that projects the input $Y$ into a higher-dimensional target space, and $R(\cdot)$ is a learnable data-driven reconstruction module, such as CNN-based neural networks. A residual connection is employed to facilitate learning and enhance reconstruction quality.
@@
However, this paradigm faces inherent limitations when applied to spectral reconstruction. For the upsampler $U(\cdot)$, commonly used techniques such as bilinear interpolation and transposed convolution are less effective in spectral super-resolution, as the upsampling ratio is substantially higher than in typical spatial super-resolution tasks. Additionally, this paradigm is entirely data-driven, lacking integration of physical processes. Besides, most data-driven reconstruction modules rely on deep neural networks that learn mappings between finite-dimensional spaces, neglecting the continuity of spectral signals, thereby limiting both performance and applicability.
@@
Hence, we adopt a similar yet more effective three-stage design, as illustrated in Figure 1. Given the MSI $Y$, wavelength coordinates $\boldsymbol{w}$, and the SRF $S_{\boldsymbol{w}}$ discretized at $\boldsymbol{w}$, we sequentially apply the following three stages to obtain the final HSI estimate.
@@
    Upsampling: In the first upsampling stage, we apply the proposed ACP method to obtain the initial upsampled HSI, which can be formulated as
    $\bar{X}=U(Y|Z_{\boldsymbol{w}},S_{\boldsymbol{w}})=P_{\text{ACP}}(Z_{\boldsymbol{w}},S_{\boldsymbol{w}},Y),$ (6)
    where $U(\cdot|Z_{\boldsymbol{w}},S_{\boldsymbol{w}})$ is the conditional upsampler guided by $Z_{\boldsymbol{w}}$ and $S_{\boldsymbol{w}}$, $P_{\text{ACP}}(\cdot,\cdot,\cdot)$ is our proposed angular-consistent projector with given pixel-wise prior $Z_{\boldsymbol{w}}\in\mathbb{R}^{C\times N}$ as the angular guidance, SRF $S_{\boldsymbol{w}}$ and MSI $Y$, which are all discretized at $\boldsymbol{w}$.
@@
    $\tilde{X}=R_{no}(\bar{X},\boldsymbol{w})+\bar{X},$ (7)
    where $\tilde{X}$ is the reconstructed HSI, $R_{no}(\cdot,\cdot)$ is the neural operator that takes the functional value and the coordinate as inputs, and a residual connection is employed similar to Equation 5.
@@
    Refinement: In the final refinement stage, we project the output to the solution space of (4), which imposes a hard constraint on our result. This stage can be formulated as
    $\hat{X}=R_{f}(\tilde{X}|S_{\boldsymbol{w}},Y)=P_{\text{ACP}}(\tilde{X},S_{\boldsymbol{w}},Y),$ (8)
    where $\hat{X}$ is the output HSI, $R_{f}(\cdot|S_{\boldsymbol{w}},Y)$ is the refinement operator. Although the $R_{f}(\cdot|S_{\boldsymbol{w}},Y)$ employs the same ACP procedure as the upsampler in the first stage, its objective is fundamentally different. The upsampling stage aims to produce a physically plausible upsampled HSI. In contrast, the refinement stage projects the output onto the solution space of (4), thereby enforcing the corresponding hard constraint.
Our framework successfully addresses the limitations of the traditional two-stage paradigm. First, the ACP method embeds physical priors into the data-driven framework through pixel-wise guidance. It also refines the output by solving a constrained optimization problem, thereby enforcing the hard constraint defined in (4). Moreover, previous studies have shown that neural operators can accommodate inputs defined on meshes with different discretizations [22, 18], so we employ a neural operator backbone to address the inherent limitation of conventional neural networks that can only model discrete spectral vectors.
Within the proposed framework, the angular-consistent projection method and the reconstruction neural operator serve as two core components, which are detailed in Section III-B and Section III-C, respectively.
III-B Angular-Consistent Projection
In this section, we propose a novel method termed Angular-Consistent Projection, based on the formulation in Section III-A. We seek a physically plausible HSI estimate that remains close to the prior $Z$ while satisfying the constraint in (4). To achieve this, we can consider the following affine space projection problem:
$\min_{X}$ $\quad\text{SAM}(X,Z)$ (9)
s.t. $\quad Y=SX,$